Directions: Show your complete solutions legibly for every item




1. Given the following reaction:




Mg(OH)2+ 2 HCI - MgCl2+ 2 H20




If 14.5.0 g of Mg(OH)2 and 12.0 g of HCI are combined




a. What is the limiting reactant?




b. How many grams of Mg Cl2 will be produced?

1
Expert's answer
2022-03-30T14:12:02-0400

M(Mg(OH)2) = 58 g/mol;

m(Mg(OH)2) = 14.5 g;

M(HCl) = 36.5 g/mol;

m(HCl) = 12.0 g;

M(MgCl2) = 95 g/mol;

n(Mg(OH)2) = m(Mg(OH)2)/M(Mg(OH)2) = 14.5/58 = 0.250 mol;

n(HCl) = m(HCl)/M(HCl) = 12.0/36.5 = 0.329 mol;

Mg(OH)2+ 2HCI = MgCl2+ 2H2O

By the chemical reaction:

n(Mg(OH)2)' = n(Mg(OH)2) = 0.250 mol;

n(HCl)' = 1/2 * n(HCl) = 1/2 * 0.329 = 0.164 mol.

So, HCl is the limiting reactant and Mg(OH)2 is in excess.

n(MgCl2) = 1/2 * n(HCl) = 1/2 * 0.329 = 0.164 mol.

m(MgCl2) = n(MgCl2) * M(MgCl2) = 0.164 * 95 = 15.6 g.

Answer: HCl is the limiting reactant; m(MgCl2) = 15.6 g.